Directions: Solve the following problems using GRESA and answer the last question.

1. Calculate the pressure exerted by a 0.25 mole sulfur hexafluoride in a steel vessel 

having a capacity of 1 250 mL at 70.0 0C.

2. Fermentation of glucose produces gas in the form of carbon dioxide. How many moles 

of carbon dioxide is produced if 0.78 L of carbon dioxide at 20.1 0C and 1.00 atm was 

collected during the process? 

3. How was the ideal gas equation derived? Explain.



1
Expert's answer
2021-06-10T03:34:10-0400

1. n=0.25 mol

V=1250 mL = 1.25 L

T=70.0 + 273=343 K

Ideal Gas Law:

pV=nRT

R = 0.08206 L×atm/mol×K

"p= \frac{nRT}{V} \\

= \frac{0.25 \times 0.08206 \times 343}{1.25} \\

= 5.63 \;atm"

Answer: 5.63 atm


2. V=0.78 L

T=20.1 + 273 = 293.1 K

p=1 atm

"n = \frac{pV}{RT} \\

= \frac{1 \times 0.78}{0.08206 \times 293.1} \\

= 0.032 \;mol"

Answer: 0.032 mol


3. PV = nRT

P is the pressure of the ideal gas.

V is the volume of the ideal gas.

n is the amount of ideal gas measured in terms of moles.

R is the universal gas constant.

T is the temperature.